OpenStudy (anonymous):

Find the lowest common denominator that would eliminate the fractions in this equation. 5x/6-x/3+1=8/9

OpenStudy (anonymous):

18

OpenStudy (anonymous):

is the answer just simplly 18 @myko

OpenStudy (anonymous):

is that the answer? 18

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[5 x/6 - x/3 + 1 = 8/9 \to 1+\frac{x}{2}= \frac{8}{9}\]1*2*3*3=18 Multiply each side by 18.\[18\left(1+\frac{x}{2}\right)=18\left(\frac{8}{9}\right) \]\[18+9 x= 16\]\[x=-\frac{2}{9} \]